OpenStudy (rayman5766):

What is the solution to the equation? 7(5 + 9) = 7 · x + 7 · 9 A. x = 1 B. x = 5 C. x = 7 D. x = 9

OpenStudy (anonymous):

\[7\left( a+b \right)=7*a+7*b\]

OpenStudy (rayman5766):

@jordanloveangel

OpenStudy (jordanloveangel):

http://www.themathpage.com/alg/absolute-value.htm best i can do